OpenStudy (anonymous):

Convert 45°C to Fahrenheit.

OpenStudy (bibby):

Converting celcius to Fahrenheit = Multiply by 9, then divide by 5, then add 32

OpenStudy (anonymous):

Thanks(:

OpenStudy (bibby):

might as well finish this in the question.\[(\frac{ (45) }{ }* \frac{ 9 }{ 5 })+32 = (9*9)+32 = 81+32 = 113\]

OpenStudy (anonymous):

yea thats exactly what I got thanks